# Mad Libs Are All You Need: Augmenting Cross-Domain Document-Level Event Argument Data

### v1

@@ 4.2 Baseline EAE Models
For each experiment, we first fine-tune the model on the source-domain training set, then on the few-shot target dataset. Target domain models are chosen using validation set performance. We train all models on Google Colab using an NVIDIA A100 GPU. Throughout the course of the project we estimate an upper-limit of 75 compute hours. Additional information regarding model hyperparameters, compute resources, and training details can be found in Appendix A.
@@ 4.3 Baseline Augmentation Methods
We compare the performance of MLA with a set of suitable augmentation baseline methods. Note that MLA is the first of its kind to generate document-level event arguments. Thus, as baselines we adapt two methods (i) Mask-then-Fill Gao et al. (2022): An augmentation method developed for sentence-level event extraction. This experiment will evaluate how prior models on sentence-level tasks will perform on document-level texts. (ii) Struct2Text: This is a custom baseline we develop which maps event records to documents using LLMs. We base our implementation on a synthetic data generation pipeline provided by LangChain. 444python.langchain.com
  • Mask-then-Fill (MtF): A sentence-level event extraction augmentation strategy which first masks contiguous spans of non-annotated text, then uses T5 Raffel et al. (2020) to fill in the missing spans. Following the implementation of Gao et al. (2022), we first fine-tune T5 on the Gigaword corpus Graff et al. (2003) to familiarize the model with news texts. Next, we split a given training document into event-related and non-event-related fragments, where an event-related fragment is any annotated span. We then mask one non-event-related fragment, and use the fine-tuned T5 model to fill in the masked context.
  • Struct2Text (S2T): We use the data augmentation tool provided by LangChain to (i) generate structured event records in a few-shot setting, then (ii) convert the new event record into a document. The document should contain the exact argument mentioned in the newly generated event record. However, in the case where it fails to do so, we take the closest n-gram in the generated text which matches the argument span in the event record as the annotated span. We use SBERT 555sentence-transformers/all-MiniLM-L6-v2 to encode the spans, and cosine similarity to find the closest matching n-gram. More details of this baseline can be found in Appendix B.3.
For Mad Lib Aug and Struct2Text, we use GPT-3.5 for all text generation. We choose GPT-3.5 due to its capacity to reliably model JSON data, and its relative affordability when compared to GPT-4. All models use default OpenAI parameters with temperature = 1.0. Please refer to Appendix B for additional details on this experiment.
@@ 4.4 Evaluation Metrics
We evaluate each model using an exact match F1 score as motivated by prior work Tong et al. (2022). Additionally, we evaluate on a diverse suite of EAE metrics which aim to capture performance on challenging role types. Inspired by Li et al. (2022), for each of the following metrics, we subset samples based on role-type and compute mean F1 score per-role .
  • Role-F1: This metric removes bias towards prediction of common roles. We compute the F1 score of each role type, then report the macro average across all roles. There are 60 total roles in the target domain.
@@
  • Few-Shot F1: This metric identifies performance on roles with low representation in the DocEE target training set. We compute F1 score of each role with less than 5 target domain training samples, then report macro average. There are 31 roles in this subset.
  • RDF1: Our novel metric, RDF1, computes F1 score of each role in the bucket of semantic outliers, then reports the macro average. Statistical depth identifies 13 roles as outliers which are used in this evaluation. See section 3.2 for additional details.
For each metric, we consider the exact match between a set of predicted and ground truth tuples consisting of (document id, role type, argument). We report the mean and standard deviation of three experimental trails with different random seeds for each metric.
@@ 5 Results
MLA Boosts Overall Performance:
Our results are shown in Table LABEL:tab:results. We find that for both BERT-QA and LongFormer-Seq, Mad Lib Aug increases overall F1 score by an average of 2.6 F1 points. Interestingly, for BERT-QA we find that training with only synthetic Mad Lib augmentation data, we out-perform the baseline which trains using only real documents across all metrics. This result follows from MLA’s ability to generate high-quality data while maintaining a more even distribution of roles when compared to the raw training data.
We find similar results on the Role-F1 metric, where all models receive a significant boost in performance by employing generative data augmentation. While our results show its advantageous to employ any form of data augmentation to improve Role-F1, MLA stands out as the highest performing model as it increases performance by an average of 4.4 F1 points. Notably, MLA augmentation also has much lower variation across multiple trials when compared to Mask-then-Fill and Struct2Text augmentation on BERT-QA.
MLA Boosts Performance on 0 and Few-Shot Roles:
Our results highlight the significant increase in performance data augmentation provides to 0 and few-shot roles. While both MLA and Struct2Text improve baseline performance on these metrics, MLA is the superior augmentation method on average. When focusing on the Real + Aug MLA experiment, we see that MLA improves baseline performance by an average of 3.9 points and 5.2 points for 0-shot and few-shot roles, respectively. Experiments using Mask-then-Fill augmentation perform on-par or worse than those using no augmentation. This is intuitive as Mask-then-Fill is augmenting existing data, thus it is unable to provide additional information about 0 and few-shot roles.
@@
Our results highlight that all of our generative data augmentation baselines significantly improve performance on the Role-Depth F1 metric. On BERT-QA, Struct2Text out-performs MLA but with significant variability between experiments. Overall, both methods provide a significant increase in RDF1, which was the most challenging subset for the BERT-QA Real Only baseline. On LongFormer-Seq, MLA out-performs Struct2Text by 0.6 RDF1. Both Struct2Text and MLA provide a 5.4 and 6.0 boost over our no augmentation baseline respectively.
@@ 6 Conclusion
In this study, we showcase the capacity of LLMs to generate event-centric document-level argument extraction data. Our method, Mad Lib Aug, produces high-quality DocEAE annotation and significantly improves performance on a variety of cross-domain DocEAE metrics. Mad Lib Aug focuses on generating large categorically templated documents. This method circumvents issues surrounding LLMs failing to follow data generation instructions/constraints, such as including/excluding certain roles in the text. Another key feature of Mad Lib Aug is that it is relatively cost-effective with respect to the number of LLM iterations needed to produce the final output. A future non-template generation approach may require significant self-reflection and/or iteration before producing a document with accurate EAE augmentation. In the future, we hope to explore how Mad Lib Aug can generate documents with greater diversity, from both the Mad Lib Generation perspective (e.g. generating narrative works) and the Mad Lib Slot-Filling perspective (e.g. the Mad Lib solver takes the perspective of a specific person). Additional future work may aim to extend this method to work with trigger-centric event datasets.
@@ 7 Limitations
A limitation of our work is use of a closed-source model, GPT-3.5. However, OpenAI models are well-suited to our approach due to their significant capacity to model JSON data  [^6] https://platform.openai.com/docs/api-reference/chat. Future works may explore how to adapt smaller, open-source LLMs for this task as well as the benefit of employing higher-performing models such as GPT-4.
Another limitation of this work is that we did not do an ablation study exploring the following. (i) How varying the number and selection of in-context demonstrations for LLM impacted the overall performance. (ii) How varying the number $n$ of augmented documents per-event type impacted performance. However, we believe that the results in Table LABEL:tab:results clearly demonstrate that MLA produces high-quality DocEAE samples. Future works can explore how to scale this method in effort to further increase the overall performance.
This work is also limited in that we do not explore the capacity of our model to generate large-scale data augmentation with fine-grained control over the role distribution. This was deemed out of scope for this work due to the cost of data generation and the focus on a few-shot task. Additionally, this is an extremely difficult sub-task which requires a more targeted solution. Future works will explore scaling our framework in this way, on both in and cross-domain formatted tasks.
@@ Appendix A Baseline Models
We will make all models available on the Huggingface to ensure reproducibility. Details of our baseline models are found in the following sections.
@@ A.1 BERT-QA
All of the BERT-QA experiments are run using the Huggingface Wolf et al. (2019) Transformers library. We use a BERT-base model (110m parameters) fine-tuned on the SQUAD-v2 dataset 777deepset/bert-base-uncased-squad2 as our base model. Given that BERT has a max token length of 512, we process longer documents in chunks using a stride of 50 tokens. We then fine-tune the model for 3 epochs on the source domain using default Huggingface Trainer parameters  [^8] https://huggingface.co/docs/transformers/main_classes/trainer. During cross-domain fine-tuning, we again train for 3 epochs on target domain data with default parameters and use the model with the best validation loss score for test set prediction. During inference, we take up to 2 predictions per-question, as some documents express more than one argument per role. We use a batch size of 48 for all BERT-QA experiments and train our models on Google Collab use an NVIDIA A100 GPU.
@@ A.2 LongFormer-Seq
All LongFormer-Seq experiments use the LongFormer-base (149M parameters) 999allenai/longformer-base-4096 and are run using the Huggingface Transformers library. As with BERT-QA, we fine-tune the model for 3-epochs on the source training set with default parameters. During cross-domain fine-tuning, we fine-tune for a maximum of 10 epochs using validation F1 score from SeqEval Nakayama (2018) to determine the best model. Note we train this model longer as a new classification head must be learned for cross-domain learning. We use a batch size of 4 for all LongFormer experiments and train our models on Google Collab use an NVIDIA A100 GPU.
@@ B.1 Hyperparameters
For each model, we use “gpt-3.5-turbo-0125" from OpenAI. We use temperature=1 with all other parameters left default. We use k=5 in-context learning examples when applicable.
B.2 Mad Lib Aug
B.2.1 Design Choices
Semantic Category Matching:
Mad Lib Aug (MLA) provides a way to generate templates for EAE data, where masked phrases can be string-replaced, removing the problem of reliably generating arguments via LLM instruction. However, given that we are still generating long templates, there is still the issue of reliably generating the masked categories. We find that occasionally, the model will generate a slightly different category, e.g. removing the plurality of the category or adding an additional word. We resolve this by simply computing the cosine similarity (as determined by SBERT 101010sentence-transformers/all-MiniLM-L6-v2) between the all categories and the generated category and returning the best match. We observe that this method reliably returns the correct category.
Few-Shot Sampling Strategy:
When random sampling from the source domain for few-shot Mad Lib generation, we randomly sample from the subset of source domain samples with character length between $\approx 1000-1500$ to manage API costs. From that subset, we focus on sampling from the 20% of documents which are more role-dense.
B.2.2 Prompts
@@
You are a creative Mad Lib writer.
Write a madlib for a “[event type]" event in the style of a newspaper article. The number of categories in the text should be exactly [num categories]. A category can occur more than once in the text. Use the following categories for Mad Lib blanks:
Categories: [categories]
@@
Fill-in missing information in the document about a "[event type]" event.
For each bracketed [CATEGORY] in the document, fill-in the missing information and add it to a JSON object. Assume all information about the event is already known and publically available.
Make sure the ANSWER written for each [CATEGORY] fits seemlessly into the document. You should output a JSON of general format "CATEGORY: ANSWER".
Document: [document]
JSON:
B.3 Struct2Text
Our Struct2Text baseline follows the synthetic data generation tutorial provided by LangChain 111111python.langchain.com/docs/use_cases/data_generation. This method is based on PyDantic  [^12] https://github.com/pydantic/pydantic and OpenAI Function Calls.
The strategy is to first few-shot generate structured event records. Specifically, we first randomly generate structures as strings from the source domain of the form “role:argument, role:argument, …". We then few-shot prompt the LLM with example structures, while providing it with a PyDantic event schema for our target domain class. The model then generates an argument for each role in the event schema. Finally, the generated event structure is fed to the LangChain OpenAI data generator. Note that we slightly adapt the prompt provided by LangChain in this method to promote generation of documents instead of sentences.
At the sentence level, we observe this method works relatively well with respect to generating texts which follow the event structure (i.e. producing the exact argument found in the event structure). Unfortunately, for long documents, larger schemas, and complex arguments, the LLM tends to deviate away from the structure template. Since we need exact mentions to be found in the text for extractive EAE, we perform semantic n-gram matching with SBERT-encoded n-grams, which identifies the n-gram in the generated document with the highest cosine similarity to the argument in the generated event schema. We empirically choose a threshold of 0.8 similarity for an n-gram to be considered a good match. If the best match is less than 0.8 similarity, we assume the model has not generated any well-matching span and thus do not include annotation for this argument.
We note that on rare occasion, the model can produce structures that fail to parse into the PyDantic framework. In such cases, we simply re-run the model and let the LLM try again. This often quickly resolves itself.
Appendix C Mad Lib Aug Example Data
C.1 Example 1
Generated Mad Lib
[PEOPLE/ORGANIZATION WHO PREDICTED THE DISASTER] successfully forecasted the arrival of [STORM NAME], which hit [STORM HIT LOCATION] on [STORM HIT TIME]. This massive storm, which formed in [STORM FORMATION LOCATION] on [STORM FORMATION TIME], brought torrential rain and wind speeds of up to [MAXIMUM WIND SPEED] mph.
residents were urged to take cover as the storm approached. The storm warning level was raised to [STORM WARNING LEVEL], indicating the severity of the weather system. Many areas experienced [AMOUNT OF PRECIPITATION] of rain, leading to flooding in low-lying areas and road closures.
The storm’s movement speed was recorded at [STORM MOVEMENT SPEED] mph, with a direction that headed towards [STORM DIRECTION]. The storm center passed over [STORM CENTER LOCATION], causing power outages and property damage in its wake.
Despite the efforts of emergency response teams, the storm’s impact left many without power and access to essential services. Residents were advised to stay indoors until the storm passed to ensure their safety. The influence of [STORM NAME] was felt by all in its path, prompting communities to come together in the aftermath to support one another in the recovery process.
Generated JSON Slot Filling
{’PEOPLE/ORGANIZATION WHO
PREDICTED THE DISASTER’: ’The National
Weather Service’, ’STORM NAME’:
’Hurricane Florence’, ’STORM HIT
LOCATION’: ’North Carolina’,
’STORM HIT TIME’: ’September 14, 2018’,
’STORM FORMATION LOCATION’:
’the Atlantic Ocean’,
’STORM FORMATION TIME’:
’September 4, 2018’,
’MAXIMUM WIND SPEED’: ’130’,
’STORM WARNING LEVEL’: ’Category 4’,
’AMOUNT OF PRECIPITATION’:
’over 30 inches’,
’STORM MOVEMENT SPEED’: ’5’,
’STORM DIRECTION’: ’northwest’,
’STORM CENTER LOCATION’: ’Wilmington’}
Final Document
The National Weather Service successfully forecasted the arrival of Hurricane Florence, which hit North Carolina on September 14, 2018. This massive storm, which formed in the Atlantic Ocean on September 4, 2018, brought torrential rain and wind speeds of up to 130 mph.
North Carolina residents were urged to take cover as the storm approached. The storm warning level was raised to Category 4, indicating the severity of the weather system. Many areas experienced over 30 inches of rain, leading to flooding in low-lying areas and road closures.
The storm’s movement speed was recorded at 5 mph, with a direction that headed towards northwest. The storm center passed over Wilmington, causing power outages and property damage in its wake.
Despite the efforts of emergency response teams, the storm’s impact left many without power and access to essential services. Residents were advised to stay indoors until the storm passed to ensure their safety. The influence of Hurricane Florence was felt by all in its path, prompting communities to come together in the aftermath to support one another in the recovery process.
C.2 Example 2
Generated Mad Lib
A devastating earthquake of magnitude [MAGNITUDE] struck the region of [AFFECTED AREA] on [DATE], resulting in catastrophic damage and significant loss of life. The earthquake’s epicenter was located near [EPICENTER], causing widespread destruction across the region.
In the aftermath of the earthquake, [AID AGENCY] quickly mobilized to provide emergency assistance to those affected, delivering [AID SUPPLIES/AMOUNT] in relief supplies to the affected areas. The aid agency also set up temporary settlements to accommodate displaced residents and provide them with basic necessities.
The earthquake led to the destruction of [NUMBER OF DESTROYED BUILDING] buildings, leaving many people homeless. Efforts are underway to rebuild [NUMBER OF REBUILDING HOUSE] homes for those who lost everything in the disaster. Rescue teams worked tirelessly to save [NUMBER OF TRAPPED PEOPLE] people trapped under the debris, managing to rescue [NUMBER OF RESCUED PEOPLE] individuals.
Unfortunately, the earthquake also resulted in a significant number of casualties, with [CASUALTIES AND LOSSES] people losing their lives in the disaster. Many more were injured and had to be evacuated to safety. The economic loss caused by the earthquake is estimated to be around [ECONOMIC LOSS], as the region faces a long road to recovery.
The aftershocks following the earthquake have been frequent, with [NUMBER OF AFTERSHOCKS] tremors further destabilizing the region. The affected area continues to be monitored closely for any potential risks, as residents remain on high alert for any further seismic activity. Our thoughts and prayers are with all those affected by this tragic event.
Generated JSON Slot Filling
{’MAGNITUDE’: ’7.2’,
’AFFECTED AREA’: ’Central Asia’,
’DATE’: ’March 12, 2022’,
’EPICENTER’: ’the town of
Shakhimardan’,
’AID AGENCY’: ’International Red Cross’,
’AID SUPPLIES/AMOUNT’:
’over 100,000 tons’,
’NUMBER OF DESTROYED BUILDING’:
’over 2,000’,
’NUMBER OF REBUILDING HOUSE’: ’1,500’,
’NUMBER OF TRAPPED PEOPLE’: ’50’,
’NUMBER OF RESCUED PEOPLE’: ’38’,
’CASUALTIES AND LOSSES’: ’at least 250’,
’ECONOMIC LOSS’: ’\$50 million’,
’NUMBER OF AFTERSHOCKS’: ’over 100’}
Final Document
A devastating earthquake of magnitude 7.2 struck the region of Central Asia on March 12, 2022, resulting in catastrophic damage and significant loss of life. The earthquake’s epicenter was located near the town of Shakhimardan, causing widespread destruction across the region.
In the aftermath of the earthquake, International Red Cross quickly mobilized to provide emergency assistance to those affected, delivering over 100,000 tons in relief supplies to the affected areas. The aid agency also set up temporary settlements to accommodate displaced residents and provide them with basic necessities.
The earthquake led to the destruction of over 2,000 buildings, leaving many people homeless. Efforts are underway to rebuild 1,500 homes for those who lost everything in the disaster. Rescue teams worked tirelessly to save 50 people trapped under the debris, managing to rescue 38 individuals.
Unfortunately, the earthquake also resulted in a significant number of casualties, with at least 250 people losing their lives in the disaster. Many more were injured and had to be evacuated to safety. The economic loss caused by the earthquake is estimated to be around \$50 million, as the region faces a long road to recovery.
The aftershocks following the earthquake have been frequent, with over 100 tremors further destabilizing the region. The affected area continues to be monitored closely for any potential risks, as residents remain on high alert for any further seismic activity. Our thoughts and prayers are with all those affected by this tragic event.

### v2

@@ 4.2 EAE Models
For each experiment, we first fine-tune the model on the source-domain training set, then on the few-shot target dataset. Target domain models are chosen using validation set performance. We train all models on Google Colab using an NVIDIA A100 GPU. Throughout the course of the project we estimate an upper-limit of 100 compute hours. Additional information regarding model hyperparameters, compute resources, and training details can be found in Appendix A.
@@ 4.3 Baseline Augmentation Methods
While this work is the first to explore document-level synthetic data generation, we compare our approach to a relevant EAE augmentation method developed for a sentence-level task and can be adapted for DocEE: Mask-then-Fill (MTF) Gao et al. (2022). MTF augments DocEE documents by masking a single contiguous span of non-annotated text in each document and infilling the mask using a T5 Raffel et al. (2020) model fine-tuned on the Gigaword corpus Graff et al. (2003). Since MTF was developed for sentence-level tasks, we additionally implement a document-level variant: Doc-MTF. Doc-MTF is similar to MTF, except the T5 model is trained to fill masks of larger contiguous spans of non-annotated text in documents from the MultiNews dataset Fabbri et al. (2019), which is more similar in style and length to DocEE texts than the original Gigaword corpus. Additionally, Doc-MTF is trained to fill multiple masks in a single document, which enables more diverse augmentations than the original single-mask variant.
4.4 LLMs for Data Generation
Since MLA uses an in-context learning-based prompting strategy, we can leverage both small open-source LLMs as well as high-powered proprietary LLMs. We thus evaluate MLA in the context of (i) Llama2-7b, (ii) Llama2-70b (iii) GPT-3.5 (iv) GPT-4. S2T, on the contrary, relies on function calling techniques which OpenAI models are specifically fine-tuned to utilize, making it challenging to integrate with smaller open-source LLMs. We thus only investigate S2T in the context of GPT-3.5 and GPT-4. We use temperature = 1 for each LLM experiment. Please refer to Appendix B for additional details on LLMs used in this study.
@@ 4.5 Evaluation Metrics
We evaluate each model using an exact match F1 score as motivated by prior work Tong et al. (2022). Additionally, we evaluate on a diverse suite of EAE metrics which aim to capture performance on challenging role types. Inspired by Li et al. (2022), for each of the following metrics, we subset evaluation data based on role-type and compute mean F1 score per-role.
  • Role-F1: We compute the F1 score of each role type, then report the macro average across all roles. This metric removes bias towards prediction of common roles. Role bias is removed in this setting as high-frequency roles no longer disproportionately affect overall F1 score. There are 60 total roles in the target domain.
@@
  • RDF1: Our novel metric, RDF1, computes the F1 score of each role in the bucket of semantic outliers, then reports the macro average. Statistical depth identifies 13 roles as outliers which are used in this evaluation. See section 3.3 for additional details. Note that 92% of these roles are few-shot, meaning they have less than 5 training examples.
[…]
@@ 5 Results
Significant Improvement in F1 Score on Zero-Shot Roles:
We find that both MLA and S2T have the capacity to significantly improve the breadth of roles which can be accurately predicted by a DocEAE model. On BERT-QA, MLA and S2T provide a 13-pt and 14-pt boost in Z-Shot F1 score respectively. The LF-Seq model also achieves a 14-pt boost in F1 score for MLA and S2T when compared to baseline models. We note that there is little difference between the results using GPT-3.5 and Llama2-70b across all experiments, highlighting the capacity of MLA to work with open-source LLMs. In general, we find that BERT-QA out-performs LF-Seq in all settings, which is intuitive as BERT-QA is only focused on one role per-inference, as where LF-Seq has the potential to accidentally assign labels from other event types.
@@
Both MLA and S2T significantly improve performance on semantically outlying RDF1 roles. Notably, our smallest LLM, Llama2-7b, which has less representational capacity compared to other LLMs still generates samples with highly useful training signal. MLA and S2T improve performance over the nearest baseline on RDF1 by 11-pts and 8-pts respectively on the BERT-QA model. On the LF-Seq model, we see a 9-pt and 10-pt boost in RDF1 from MLA and S2T respectively. This result confirms that LLM-based DocEAE generation significantly enhances cross-domain learning. We note that for all metrics, MLA and S2T models achieve similar F1 scores, highlighting how this method leads to more evenly distributed role performance vs baseline models which are skewed to overall F1.
Trade-Off Between Overall F1 and Role-Specific Metrics:
Our results show that across all experiments there is no statistically significant difference in overall F1 score between baseline and augmented models. We find that roles which have significant representation in both the training and testing set often do not benefit from augmentation. This is intuitive, as the model can overfit to high training frequency roles and improve overall F1 since the test set is dominated by a small set of common roles. For example, the role Date, which appears 28 times in the training set (3rd most frequently occurring training role), and 1841 times in the test set, does not benefit from augmentation. A model can thus overfit to Date and gain significant performance increase in overall F1 score as this role type dominates the test set. Additionally, it may be hard to generate diverse augmentations for date, as DocEE considers arguments such as “Now" or “just as the lunch hour was ending Tuesday" to be dates, while MLA GPT-4 exclusively generates things in some variation of the Day-Month-Year format. This is in contrast to the role The State of the Volcano (Dormant or Active) which occurs only 52 times in the test set, and never occurs in the training set. In this situation, the LLM significantly improves model performance as it has no prior exposure to such a role-type. Thus, if it is more important to the end user of a DocEAE system that a wide variety of roles can be accurately extracted, both MLA and S2T are well-suited to improve performance in low-resource settings.
Performance Plateaus at 8x Augmentation:
In Table 2, we explore how scaling up augmentation impacts performance on three different metrics. We find that increasing the scale from 0x to 4x provides significant improvement. Additionally, scaling from 4x to 8x improves performance on Z-Shot F1 and RDF1. However, further scaling the data to 16x has limited impact and can even reduce performance on Z-Shot F1 and RDF1. We believe the limits of scaling may be driven by the 0-shot nature of the generation, where it can be challenging for an LLM to come up with diverse arguments given only event type and role name information. For example, many of the events pertaining to natural disasters have the role of “Aid Agency", for which GPT-4 overwhelmingly generates “Red Cross" or “American Red Cross" as the argument, as it is one of the most popular Aid Agencies in the world. Future works may address this gap via external knowledge injection during the data generation process.
Aug Scale F1 Z-Shot F1 RDF1
Baseline - 0.31 0.19 0.20
MLA-GPT4 4x 0.31 0.30 0.31
8x 0.31 0.32 0.32
16x 0.30 0.32 0.31
S2T-GPT4 4x 0.30 0.33 0.29
8x 0.30 0.36 0.31
16x 0.28 0.35 0.28
Table 2: Exploring how augmentation scale affects performance. We find that for Z-Shot F1 and RDF1, 8x augmentation provides the strongest performance for both models. There are thus limits to how much scaling augmentation can improve performance.
@@ 6 Conclusion
In this study, we showcase the capacity of LLMs to generate event-centric document-level argument extraction data. Our LLM-driven framework produces high-quality DocEAE annotation and significantly improves performance on a variety of cross-domain DocEAE metrics. We find that MadLib Aug is a promising method of generating categorized, templated documents. This method is shown to be successful using both open-source and proprietary models, making it an accessible form of DocEAE augmentation. S2T, on average, provides slightly higher scores than MLA on various metrics and is thus a powerful method of DocEAE data generation. In the future, we hope to explore how Mad Lib Aug can generate documents with greater diversity, from both the Mad Lib Generation perspective (e.g. generating narrative works) and the Mad Lib Slot-Filling perspective (e.g. the Mad Lib solver takes the perspective of a specific person). We also wish to explore adding role definitions to the generation process for both MLA and S2T. Additional future work may aim to extend this method to work with trigger-centric event datasets.
@@ 7 Limitations
This work is limited in that we do not explore the capacity of our model to generate large-scale data augmentation with fine-grained control over the role distribution. This was deemed out of scope for this work due to the cost of data generation and the focus on a few-shot task. Additionally, this is an extremely difficult sub-task which requires a more targeted solution. Future works will explore scaling our framework in this way, on both in and cross-domain formatted tasks.
A limitation of the S2T model is it’s reliance on the OpenAI function calling capabilities. Future works may wish to explore how to adapt this method to an open-source LLM which has been fine-tuned for function calling in a manner that facilitates DocEAE data generation. Another general limitation of this study is that we do not thoroughly explore the capacity of these methods to generate multi-argument roles. For example, if there are two different causes to an event, having the MLA model generate templates for [Cause #1] and [Cause #2]. Such extensions will be the focus of future work.
Finally, this work is limited in that it does not fully explore the capacity of LLMs to perform related types of augmentation such as style transfer (e.g. generating Mad Libs in a variety of styles). This may improve performance even further and will be a subject of future work.
@@ Appendix A Baseline Models
Details of our baseline models are found in the following sections.
@@ A.1 BERT-QA
All of the BERT-QA experiments are run using the Huggingface Wolf et al. (2019) Transformers library. We use a BERT-base model (110m parameters) 555bert-base-uncased as our base model. Given that BERT has a max token length of 512, we process longer documents in chunks using a stride of 50 tokens. We then fine-tune the model for 3 epochs on the source domain using default Huggingface Trainer parameters  [^6] https://huggingface.co/docs/transformers/main_classes/trainer. During cross-domain fine-tuning, we again train for 3 epochs on target domain data with default parameters and use the model with the best validation loss score for test set prediction. This typically converges after 1 epoch. To help encourage extraction of multi-argument roles, we only include unique arguments as training examples for QA-based extraction. During inference, we take up to 2 predictions per-question, as some documents express more than one argument per role. We perform inference using the Huggingface question-answering pipeline. We consider roles present in test set during evaluation. We use a batch size of 48 for all BERT-QA experiments and train our models on Google Colab use an NVIDIA A100 GPU.
@@ A.2 LongFormer-Seq
All LongFormer-Seq experiments use the LongFormer-base (149M parameters) 777allenai/longformer-base-4096 and are run using the Huggingface Transformers library. As with BERT-QA, we fine-tune the model for 3-epochs on the source training set with default parameters. During cross-domain fine-tuning, we fine-tune for a maximum of 10 epochs using validation F1 score from SeqEval Nakayama (2018) to determine the best model. Note we train this model longer as a new classification head must be learned for cross-domain learning. We use a batch size of 4 for all LongFormer experiments and train our models on Google Collab use an NVIDIA A100 GPU.
@@ B.1 Hyperparameters
For each model from OpenAI, we use temperature=1 with all other parameters left default. We use k=3 in-context learning examples when applicable. For open-source Llama models we restrict generation to 512 tokens. Future works may explore generation of even longer documents.
B.2 Mad Lib Aug Post Processing
To ensure high-quality data, we post-process our generated Mad Libs. An error that can occur during this process is hallucinated role types. For example, if the LLM invents a new role not found in the event schema when generating the Mad Lib, we may handle this in two ways: (i) we can use SBERT Reimers and Gurevych (2019), a popular textual embedding method, to encode the hallucinated category name and compute the cosine similarity between it and all possible role names for a given sample. We can often reliably match the generated category to the true category using this method for mild hallucinations (e.g. mapping “The place" back to “Location"). (ii) For smaller LLMs we find hallucinations can be more extreme, thus we simply remove the sentence containing the hallucinated role from the document. Since DocEE is largely comprised of news articles, simply removing the sentence is a non-destructive way to deal with hallucination as the documents are often comprised of a sequence of factual statements. In our experiments, we focus on method (ii) as it is less ambiguous and works with all models.
Additionally, the generation of a longer text will be particularly sensitive to the randomly drawn few-shot sample selection. Thus, there can be outlying generation attempts where Mad Libs are generated (i) with little to no role information (ii) with incorrectly used brackets (iii) with improperly formatted solutions. When this happens, we simply draw a new set of few-shot examples and try again. This issue is predominantly found when using Llama2-7b. For larger models such as GPT-4, this almost never happens. We restrict the model to a maximum of 5 retries during such situations.
B.3 Struct2Text Post Processing
Similar to MLA, S2T can suffer from random generation errors. Unlike MLA, S2T’s errors are largely constrained to JSON parsing issues which occurring during use of function calling. We thus similarly allow for 5 attempts at sample generation, which is shown to be enough to handle random JSON errors.
B.3.1 Prompts
@@
Write a madlib for a “{event_type}" event. Use the following categories for Mad Lib blanks. Make sure all categories appear in the MadLib. Do not generate any other categories.:
Categories: {categories}
@@
Solve the MadLib. Fill-in missing information in the document. Return your answer as a bulleted list of the format “### CATEGORY: ANSWER".
Make sure the following categories are in your solution: {categories}
Madlib: {document}
MadLib Solution:
Appendix C Exploring The Effect of RDF1 Percentage
To motivate our choice of 25% as the outlying role threshold for RDF1, we recompute the results using 4 separate thresholds in the range of ±$\pm 10\%$. We find that for the DocEE dataset, the threshold does not significantly alter performance.
15% 20% 25% 30% 35% Aug Method
0.24 0.22 0.2 0.19 0.2 Baseline
0.24 0.2 0.18 0.18 0.2 MTF
0.27 0.27 0.28 0.26 0.26 S2T
0.3 0.3 0.31 0.31 0.3 MLA
Table 3: RDF1 scores at varying thresholds for outlying role designation. We find that models are relatively stable given slight fluctuations in RDF1 threshold.
Appendix D Example Data
Figure 3: Example data for Tsunami Event using Struct-2-Text (GPT-4).
Figure 4: Example data for Earthquake event using Mad Lib Aug (GPT-4).
Figure 5: Example data for Droughts Event using Mad Lib Aug (Llama-7b). Note that sample quality suffers for smaller LLMs. However, such samples still serve as useful training signal as shown in our results.
Figure 6: Example data for Storm Event using Mad Lib Aug (Llama-70b).